# Self-Distilled Masked Auto-Encoders are Efficient Video Anomaly Detectors

### v1

@@ Abstract
We propose an efficient abnormal event detection model based on a lightweight masked auto-encoder (AE) applied at the video frame level. The novelty of the proposed model is threefold. First, we introduce an approach to weight tokens based on motion gradients, thus avoiding learning to reconstruct the static background scene. Second, we integrate a teacher decoder and a student decoder into our architecture, leveraging the discrepancy between the outputs given by the two decoders to improve anomaly detection. Third, we generate synthetic abnormal events to augment the training videos, and task the masked AE model to jointly reconstruct the original frames (without anomalies) and the corresponding pixel-level anomaly maps. Our design leads to an efficient and effective model, as demonstrated by the extensive experiments carried out on three benchmarks: Avenue, ShanghaiTech and UCSD Ped2. The empirical results show that our model achieves an excellent trade-off between speed and accuracy, obtaining competitive AUC scores, while processing 1670 FPS. Hence, our model is between 8 and 70 times faster than competing methods. We also conduct an ablation study to justify our design.
@@ 1 Introduction
Figure 1: Our masked auto-encoder for abnormal event detection based on self-distillation. At training time, some video frames are augmented with synthetic anomalies. The teacher decoder learns to reconstruct original frames (without anomalies) and predict anomaly maps. The student decoder learns to reproduce the teacher’s output. Motion gradients are aggregated at the token level and used as weights for the reconstruction loss. Red dashed lines represent steps executed only during training. Best viewed in color.
In recent years, research on abnormal event detection in video gained significant traction [1, 9, 16, 17, 25, 26, 27, 33, 36, 40, 41, 46, 48, 53, 55, 57, 58, 61, 65, 72, 74, 76, 78, 82, 85, 89, 91, 92, 93, 94], due to its utter importance in video surveillance. Despite the growing interest, video anomaly detection remains a complex tasks, owing its complexity to the fact that abnormal situations are context-dependent and do not occur very often. This makes it very difficult to collect a representative set of abnormal events for training state-of-the-art deep learning models in a fully supervised manner. To showcase the rarity and reliance on context of anomalies, we refer to the vehicle ramming attacks carried out by terrorists against pedestrians. As soon as a car is steered on the sidewalk, it becomes an abnormal event. Hence, the place where the car is driven (street versus sidewalk) determines the normal or abnormal label of the action, i.e. the label depends on context. Furthermore, there are less than 200 vehicle ramming attacks registered to date [^1] https://en.wikipedia.org/wiki/Vehicle-ramming_attack, confirming the scarcity of such events (even less are caught on video).
Since training anomaly detectors under a fully supervised setting is not possible, most studies dealing with abnormal event detection took a distinct path, proposing variations of outlier detection methods [3, 12, 13, 16, 20, 28, 35, 38, 40, 42, 44, 47, 49, 51, 52, 57, 58, 59, 60, 61, 62, 63, 64, 67, 76, 85, 87, 98, 96, 97, 99, 21]. Such methods treat abnormal event detection as an outlier detection task, where a normality model trained on normal events is applied on both normal and abnormal events during inference, labeling events deviating from the learned model as abnormal. Different from the mainstream path based on outlier detection, we propose an approach to augment each training video scene with synthetic anomalies, by superimposing temporal action segments from the synthetic UBnormal data set [1] on our real-world data sets. We thus introduce synthetic anomalies at training time, enabling our model to learn in a supervised manner. Additionally, we force our model to reconstruct the original training frames (without anomalies) to limit its ability to reconstruct anomalies, hence generating higher errors when anomalies occur.
@@
A large body of work on video anomaly detection has focused on employing auto-encoders (AEs) to address the task [5, 21, 26, 27, 28, 33, 46, 77, 80], relying on the poor reconstruction capabilities of these models on out-of-distribution data. Since training is carried out only on normal examples, it is expected for AEs to exhibit high reconstruction errors when anomalies occur. However, several researchers observed that AEs generalize too well [5, 33], being able to reconstruct anomalies with very high precision. Thus, to better leverage the reconstruction error of AEs in anomaly detection, researchers explored a few alternatives, from the use of dummy [33] or pseudo-anomalies [5, 26] to the integration of memory modules [27, 46, 57]. With the same purpose in mind, we propose to employ masked auto-encoders [29] in anomaly detection, introducing new ways to regulate their generalization capacity. Indeed, we go beyond employing the standard masked AE framework, and propose two novel changes to enhance the anomaly detection performance of our model. First, we propose to weight tokens based on the magnitude of motion gradients, raising the importance of tokens with higher motion in the reconstruction loss. This helps our model to focus on reconstructing tokens with high motion, and avoid reconstructing the background scene, which is typically static for surveillance cameras. Second, we integrate a teacher decoder and a student decoder into our masked AE architecture, where the student decoder learns to distill knowledge from the already optimized teacher. To reduce our processing time, we use a shared encoder for the teacher and student models, leading to a process known as self-distillation [95]. During the self-distillation process, the shared encoder is frozen. We leverage the discrepancy between the outputs given by the teacher and student decoders along with the reconstruction error of the teacher to boost anomaly detection performance. Our entire framework is illustrated in Figure 1.
@@
State-of-the-art deep anomaly detectors [9, 25, 26, 79] typically rely on a costly object detection method to increase precision, limiting the processing bandwidth to one video stream per GPU, at around 20-30 FPS. However, for real-world video surveillance, e.g. monitoring an entire city with hundreds or thousands of cameras, the processing costs of object-centric video anomaly detectors are simply too high, given their power consumption and that one GPU can cost around \$2,000. To this end, we turn our attention to developing a lightweight model (6 transformer blocks, 3M parameters), capable of processing around 67 video streams at 25 FPS, significantly reducing the processing costs. Different from competing models performing anomaly detection at the object [9, 17, 25, 26, 33, 79, 89] or spatio-temporal cube [20, 38, 47, 14, 34, 35, 56, 45, 49, 51, 67, 71, 97] levels, we present a model that takes whole video frames as input, which is significantly more efficient (see Figure 2).
@@
We carry out experiments on three benchmarks: Avenue [47], ShanghaiTech [49] and UCSD Ped2 [51]. The empirical results show that our method is 8 to 70 times faster than competing methods [25, 26, 65, 44, 46, 57, 27, 79, 56], while achieving comparable accuracy levels. Aside from the main results, we conduct an ablation study showing that our novel design choices are supported by empirical evidence.
@@
  • To further boost the performance of our model, we introduce a data augmentation approach based on superimposing synthetic anomalies on normal training videos, which enables the masked AE model to learn with full supervision.
@@ 2 Related Work
With respect to the level at which the anomaly detection is carried out, methods can be categorized into spatio-temporal cube-level methods [44, 52, 62, 63, 72, 90, 35, 45, 47, 51, 67, 71, 97], frame-level methods [44, 63, 62], and object-level methods [17, 18, 25, 26, 33, 79, 89, 9].
@@
To boost the performance of frame-level or cube-level methods, researchers explored the inclusion of various components, such as memory modules [46, 27, 57] or masked convolutional blocks [65]. Although integrating additional modules into the framework leads to accuracy gains, the procedure often comes with efficiency drawbacks. In contrast, our goal is to achieve a superior trade-off between performance and speed, with a higher focus on efficiency. As such, we design a lightweight masked AE based on convolutional vision transformer (CvT) blocks [83], and propose several upgrades resulting in a minimal time overhead. For example, we employ knowledge distillation to leverage the discrepancy between the teacher and the student models. However, to keep the processing time to the bare minimum, we resort to self-distillation [95] and use a shared encoder for the teacher and student networks.
Object-level methods. To reduce the number of false positive detections often observed for other methods, some recent studies [9, 33, 26, 25, 89, 79] proposed to look for anomalous objects rather than anomalous frames or cubes. Object-centric methods use the prior information from an object detector, enabling the anomaly detector to focus only on objects. This kind of framework boosts the accuracy by significant margins, currently reaching state-of-the-art performance [9, 79]. However, a considerable drawback is that the inference speed of the whole framework is directly conditioned by the object detector’s speed, which is often much lower than that of the anomaly detection network [33, 25]. Hence, the processing time is significantly limited. In contrast, we perform anomaly detection at the frame level, obtaining an inference speed that is between 32 to 70 times faster than object-centric models [9, 33, 26, 25, 89, 79].
@@
Distinct from the aforementioned studies, to our knowledge, we are the first to introduce a variant of self-distillation in anomaly detection. Self-distillation [95] attaches multiple classification heads at various depths to boost the classification performance of a neural classifier. In contrast, we integrate self-distillation into a masked AE, employing two decoders of different depth. Due to the shared encoder, we are able to leverage the reconstruction discrepancy between the teacher and the student with a minimal computational overhead.
@@ 3 Method
Overview. We introduce a lightweight teacher-student transformer-based masked AE, which employs a two-stage training pipeline. In the first stage, we optimize a teacher masked AE via a reconstruction loss that employs a novel weighting mechanism based on motion gradients. In the second stage, we optimize the last (and only) decoder block of a student masked AE, which shares most of the backbone (kept frozen) with its teacher. Lastly, we describe how to create training videos with synthetic anomalies and train the masked AEs to jointly predict the anomaly maps and overlook (not reconstruct) the anomalies from training frames.
@@
Self-distillation. Knowledge distillation has already shown its utility in anomaly detection [8, 11, 15, 25, 70]. Intuitively, since the teacher and student models are both trained on normal data, their reconstructions should be very similar for normal test samples. However, their behavior is not guaranteed to be similar on abnormal examples. Therefore, the magnitude of the teacher-student output gap (discrepancy) can serve as a means to quantify the anomaly level of a given sample. Unfortunately, this approach implies using both teacher and student models during inference, virtually splitting our processing speed in half. To slash the additional burden of using another model during inference, we propose to employ a variant of self-distillation [95] with a shared encoder and two decoders, a teacher and a student. More precisely, the student branches out from the original architecture after the first transformer block of the teacher decoder, essentially adding only one transformer block.
@@
Synthetic anomalies. As observed in other studies [5, 33], AEs tend to generalize too well to out-of-distribution data. This behavior is not desired in anomaly detection, since methods based on AEs rely on having high reconstruction errors for abnormal examples and low reconstruction errors for normal ones. To this end, we propose to augment the training videos with abnormal events. Since collecting abnormal training examples from the real-world is not possible, we resort to adding synthetic (virtual) anomalies. We leverage the recently introduced UBnormal data set [1] and its accurate pixel-level annotations to crop out abnormal events and superimpose them on our training videos. The resulting examples, some depicted in Figure 3, are used to augment the training set with extra data.
The synthetic examples help our model in three ways. First, in the reconstruction loss, we consider the original training frames (without superimposed anomalies) as the ground-truth, essentially forcing our model to overlook the anomalies. Formally, in Eq. (3), we use the patches $\left\{\boldsymbol{p}_{i}^{(t)}\right\}_{i=1}^{n}$ from the normal version of frame $\boldsymbol{x}_{t}$. Second, we add the anomaly map as an additional channel to our target image. In the anomaly map, we set normal pixels to 00 and abnormal pixels to $1$. This change implies that, in Eq. (3) and Eq. (4), all patches will have an additional channel. Third, we use the ground-truth anomaly map to enhance the weights defined in Eq. (2). The added synthetic anomalies do not necessarily yield motion gradients with high magnitude. Hence, it is possible to have low weights in Eq. (3) and Eq. (4) for patches that correspond to anomaly regions. This is not desirable if we want the model to learn to predict anomalies. To this end, we propose to add the anomaly maps and the gradients together, before computing the weights as in Eq. (2). Formally, in Eq. (1), we replace $\left\{\boldsymbol{r}_{i}^{(t)}\right\}_{i=1}^{n}$ with $\left\{\boldsymbol{r}_{i}^{(t)}+\boldsymbol{a}_{i}^{(t)}\right\}_{i=1}^{n}$, where $\left\{\boldsymbol{a}_{i}^{(t)}\right\}_{i=1}^{n}$ is the set of the patches extracted from the anomaly map.
@@
$\boldsymbol{o}_{t}=\lVert\boldsymbol{x}_{t}-\hat{\boldsymbol{x}}_{t}\rVert^{2}_{2}+\lVert\hat{\boldsymbol{x}}_{t}-\tilde{\boldsymbol{x}}_{t}\rVert^{2}_{2}.$ (5)
Following [14, 34], we apply spatio-temporal 3D filtering to smooth the anomaly volumes. To obtain the frame-level anomaly scores, we keep the maximum value from each output map $\boldsymbol{o}_{t}$ and subsequently apply another temporal Gaussian filter to smooth the values.

### v2

@@ Abstract
We propose an efficient abnormal event detection model based on a lightweight masked auto-encoder (AE) applied at the video frame level. The novelty of the proposed model is threefold. First, we introduce an approach to weight tokens based on motion gradients, thus shifting the focus from the static background scene to the foreground objects. Second, we integrate a teacher decoder and a student decoder into our architecture, leveraging the discrepancy between the outputs given by the two decoders to improve anomaly detection. Third, we generate synthetic abnormal events to augment the training videos, and task the masked AE model to jointly reconstruct the original frames (without anomalies) and the corresponding pixel-level anomaly maps. Our design leads to an efficient and effective model, as demonstrated by the extensive experiments carried out on four benchmarks: Avenue, ShanghaiTech, UBnormal and UCSD Ped2. The empirical results show that our model achieves an excellent trade-off between speed and accuracy, obtaining competitive AUC scores, while processing 1655 FPS. Hence, our model is between 8 and 70 times faster than competing methods. We also conduct an ablation study to justify our design. Our code is freely available at: https://github.com/ristea/aed-mae.
@@ 1 Introduction
Figure 1: Our masked auto-encoder for abnormal event detection based on self-distillation. At training time, some video frames are augmented with synthetic anomalies. The teacher decoder learns to reconstruct original frames (without anomalies) and predict anomaly maps. The student decoder learns to reproduce the teacher’s output. Motion gradients are aggregated at the token level and used as weights for the reconstruction loss. Red dashed lines represent steps executed only during training.
In recent years, research on abnormal event detection in video gained significant traction [1, 10, 17, 18, 26, 27, 28, 36, 38, 43, 44, 49, 52, 57, 58, 61, 62, 65, 69, 76, 78, 80, 83, 87, 90, 95, 97, 98, 99, 100], due to its utter importance in video surveillance. Despite the growing interest, video anomaly detection remains a complex task, owing its complexity to the fact that abnormal situations are context-dependent and do not occur very often. This makes it very difficult to collect a representative set of abnormal events for training state-of-the-art deep learning models in a fully supervised manner. To showcase the rarity and reliance on context of anomalies, we refer to the vehicle ramming attacks carried out by terrorists against pedestrians. As soon as a car is steered on the sidewalk, it becomes an abnormal event. Hence, the place where the car is driven (street versus sidewalk) determines the normal or abnormal label of the action, i.e. the label depends on context. Furthermore, there are less than 200 vehicle ramming attacks registered to date [^1] https://en.wikipedia.org/wiki/Vehicle-ramming_attack, confirming the scarcity of such events (even less are caught on video).
Since training anomaly detectors under a fully supervised setting is not possible, most studies dealing with abnormal event detection took a distinct path, proposing variations of outlier detection methods [3, 13, 14, 17, 21, 29, 37, 41, 43, 45, 47, 51, 53, 55, 56, 61, 62, 63, 64, 65, 67, 66, 68, 71, 80, 90, 92, 104, 102, 103, 105, 22]. Such methods treat abnormal event detection as an outlier detection task, where a normality model trained on normal events is applied on both normal and abnormal events during inference, labeling events deviating from the learned model as abnormal. Different from the mainstream path based on outlier detection, we propose an approach to augment each training video scene with synthetic anomalies, by randomly superimposing temporal action segments from the synthetic UBnormal data set [1] on our real-world data sets. We thus introduce synthetic anomalies at training time, enabling our model to learn in an open-set supervised manner. Additionally, we force our model to reconstruct the original training frames (without anomalies) to limit its ability to reconstruct anomalies, hence generating higher errors when anomalies occur.
@@
A large body of work on video anomaly detection has focused on employing auto-encoders (AEs) to address the task [5, 22, 27, 28, 29, 36, 49, 81, 85], relying on the poor reconstruction capabilities of these models on out-of-distribution data. Since training is carried out only on normal examples, it is expected for AEs to exhibit high reconstruction errors when anomalies occur. However, several researchers observed that AEs generalize too well [5, 36], being able to reconstruct anomalies with very high precision. Thus, to better leverage the reconstruction error of AEs in anomaly detection, researchers explored a few alternatives, from the use of dummy [36] or pseudo-anomalies [5, 27] to the integration of memory modules [28, 49, 61]. With the same purpose in mind, we propose to employ masked auto-encoders [30] in anomaly detection, introducing new ways to regulate their generalization capacity. Indeed, we go beyond employing the standard masked AE framework, and propose three novel changes to enhance the anomaly detection performance of our model. First, we propose to weight tokens based on the magnitude of motion gradients, raising the importance of tokens with higher motion in the reconstruction loss. This makes our model focus on reconstructing tokens with high motion, and avoid reconstructing the background scene, which is typically static for surveillance cameras. Second, we attach a classification head to discriminate between normal and pseudo-abnormal instances in the latent encoding space. Third, we integrate a teacher decoder and a student decoder into our masked AE architecture, where the student decoder learns to distill knowledge from the already optimized teacher. To reduce our processing time, we use a shared encoder for the teacher and student models, leading to a process known as self-distillation [101]. During the self-distillation process, the shared encoder is frozen. We leverage the discrepancy between the outputs given by the teacher and student decoders along with the reconstruction error of the teacher to boost anomaly detection performance. Our entire framework, integrating these components into a meticulous design, is shown in Figure 1.
@@
State-of-the-art deep anomaly detectors [10, 26, 27, 84] typically rely on a costly object detection method to increase precision, limiting the processing bandwidth to one video stream per GPU, at around 20-30 FPS. However, for real-world video surveillance, e.g. monitoring an entire city with hundreds or thousands of cameras, the processing costs of object-centric video anomaly detectors are simply too high, given their power consumption and that one GPU can cost around \$2,000. To this end, we turn our attention to developing a lightweight model (6 transformer blocks, 3M parameters), capable of processing around 66 video streams at 25 FPS, significantly reducing the processing costs. Different from competing models performing anomaly detection at the object [10, 18, 26, 27, 36, 84, 95] or spatio-temporal cube [21, 41, 51, 15, 35, 37, 60, 48, 53, 55, 71, 75, 103] levels, we present a model that takes whole video frames as input, which is significantly more efficient (see Figure 2).
@@
We carry out comprehensive experiments on four benchmarks: Avenue [51], ShanghaiTech [53], UBnormal [1] and UCSD Ped2 [55]. The empirical results show that our method is 8 to 70 times faster than competing methods [26, 27, 69, 47, 49, 61, 28, 84, 60], while achieving comparable accuracy levels. Aside from the main results, we conduct an ablation study showing that our novel design choices are supported by empirical evidence.
@@
  • To further boost the performance of our model, we introduce a data augmentation approach based on superimposing synthetic anomalies on normal training videos, which enables the masked AE model to learn with open-set supervision.
@@ 2 Related Work
With respect to the level at which the anomaly detection is carried out, methods can be categorized into spatio-temporal cube-level methods [47, 56, 67, 66, 76, 96, 37, 48, 51, 55, 71, 75, 103], frame-level methods [47, 66, 67], and object-level methods [18, 19, 26, 27, 49, 36, 84, 95, 10].
@@
To boost the performance of frame-level or cube-level methods, researchers explored the inclusion of various components, such as memory modules [28, 61] or masked convolutional blocks [69]. Although integrating additional modules into the framework leads to accuracy gains, the procedure often comes with efficiency drawbacks. In contrast, our goal is to achieve a superior trade-off between performance and speed, with a higher focus on efficiency. As such, we design a lightweight masked AE based on convolutional vision transformer (CvT) blocks [88], and propose several upgrades resulting in a minimal time overhead. For example, we employ knowledge distillation to leverage the discrepancy between the teacher and the student models. However, to keep the processing time to the bare minimum, we resort to self-distillation [101] and use a shared encoder for the teacher and student networks.
Object-level methods. To reduce the number of false positive detections often observed for other methods, some recent studies [10, 36, 27, 26, 49, 95, 84] proposed to look for anomalous objects rather than anomalous frames or cubes. Object-centric methods use the prior information from an object detector, enabling the anomaly detector to focus only on objects. This kind of framework boosts the accuracy by significant margins, currently reaching state-of-the-art performance [10, 84]. However, a considerable drawback is that the inference speed of the whole framework is directly conditioned by the object detector’s speed, which is often much lower than that of the anomaly detection network [36, 26]. Hence, the processing time is significantly limited. In contrast, we perform anomaly detection at the frame level, obtaining an inference speed that is between 32 to 70 times faster than object-centric models [10, 36, 27, 26, 49, 95, 84].
@@
Distinct from the aforementioned studies, to our knowledge, we are the first to introduce a variant of self-distillation in anomaly detection. Self-distillation [101] attaches multiple classification heads at various depths to boost the classification performance of a neural classifier. In contrast, we integrate self-distillation into a masked AE, employing two decoders of different depths. Due to the shared encoder, we are able to leverage the reconstruction discrepancy between the teacher and the student with a minimal computational overhead.
@@ 3 Method
Overview. We introduce a lightweight teacher-student transformer-based masked AE, which employs a two-stage training pipeline. In the first stage, we optimize a teacher masked AE via a reconstruction loss that employs a novel weighting mechanism based on motion gradients. In the second stage, we optimize the last (and only) decoder block of a student masked AE, which shares most of the backbone (kept frozen) with its teacher, to preserve efficiency. Next, we describe how to create training videos with synthetic anomalies and train the masked AEs to jointly predict the anomaly maps and overlook (not reconstruct) the anomalies from training frames. Lastly, we introduce a classification head to distinguish between frames with and without synthetic anomalies, which further boosts the performance of our method, with a marginal computational overhead.
@@
Self-distillation. Knowledge distillation has already shown its utility in anomaly detection [8, 12, 16, 26, 74]. Intuitively, since the teacher and student models are both trained on normal data, their reconstructions should be very similar for normal test samples. However, their behavior is not guaranteed to be similar on abnormal examples. Therefore, the magnitude of the teacher-student output gap (discrepancy) can serve as a means to quantify the anomaly level of a given sample. Unfortunately, this approach implies using both teacher and student models during inference, virtually splitting our processing speed in half. To slash the additional burden of using another model during inference, we propose to employ a novel variant of self-distillation with a shared encoder and two decoders, a teacher and a student. More precisely, the student branches out from the original architecture after the first transformer block of the teacher decoder, essentially adding only one transformer block.
@@
Synthetic anomalies. As observed in other studies [5, 36], AEs tend to generalize too well to out-of-distribution data. This behavior is not desired in anomaly detection, since methods based on AEs rely on having high reconstruction errors for abnormal examples and low reconstruction errors for normal ones. To this end, we propose to augment the training videos with abnormal events. Since collecting abnormal training examples from the real-world is not possible, we resort to adding synthetic (virtual) anomalies. We leverage the recently introduced UBnormal data set [1] and its accurate pixel-level annotations to crop out abnormal events and blend them in our training videos, while ensuring the temporal consistency of the added events. The resulting examples, some depicted in Figure 3, are used to augment the training set with extra data.
The synthetic examples help our model in three ways. First, in the reconstruction loss, we consider the original training frames (without superimposed anomalies) as the ground-truth, essentially forcing our model to overlook the anomalies. Formally, in Eq. (3), we use the patches $\left\{\boldsymbol{p}_{i}^{(t)}\right\}_{i=1}^{n}$ from the normal version of frame $\boldsymbol{x}_{t}$. Second, we add the anomaly map as an additional channel to our target image. In the anomaly map, we set normal pixels to 00 and abnormal pixels to $1$. This change implies that, in Eq. (3) and Eq. (4), all patches will have an additional channel. Third, we use the ground-truth anomaly map to enhance the weights defined in Eq. (2). The added synthetic anomalies do not necessarily yield motion gradients with high magnitude. Hence, it is possible to have low weights in Eq. (3) and Eq. (4) for patches that correspond to anomaly regions. This is not desirable if we want the model to detect anomalies. To this end, we propose to add the anomaly maps and the gradients together, before computing the weights as in Eq. (2). Formally, in Eq. (1), we replace $\left\{\boldsymbol{r}_{i}^{(t)}\right\}_{i=1}^{n}$ with $\left\{\boldsymbol{r}_{i}^{(t)}+\boldsymbol{a}_{i}^{(t)}\right\}_{i=1}^{n}$, where $\left\{\boldsymbol{a}_{i}^{(t)}\right\}_{i=1}^{n}$ is the set of the patches extracted from the anomaly map.
Classification head. We further harness the synthetic anomalies to train a classification head applied on the final [CLS] token of the shared encoder. The head is trained to discriminate between frames with and without synthetic anomalies. This head is trained using binary cross-entropy:
$\begin{split}\mathcal{L}_{\mbox{\scriptsize{CE}}}(\hat{\boldsymbol{x}}_{t},\theta_{E})=-y_{t}\!\cdot\!\log(\hat{y}_{t})-(1\!-\!y_{t})\!\cdot\!\log(1-\hat{y}_{t}),\end{split}$ (5)
where $y_{t}\in\{0,1\}$ is $1$ if the frame contains an anomaly and 00 otherwise, $\hat{y}_{t}$ is the prediction, and $\theta_{E}$ represents the set of weights of the shared encoder.
@@
$\boldsymbol{o}_{t}=\alpha\!\cdot\!\lVert\boldsymbol{x}_{t}-\hat{\boldsymbol{x}}_{t}\rVert^{2}_{2}+\beta\!\cdot\!\lVert\hat{\boldsymbol{x}}_{t}-\tilde{\boldsymbol{x}}_{t}\rVert^{2}_{2}+\gamma\!\cdot\!\hat{y}_{t},$ (6)
where $\alpha$, $\beta$ and $\gamma$ are hyperparameters that control the importance of the individual anomaly score components. Following [15, 35], we apply spatio-temporal 3D filtering to smooth the anomaly volumes. To obtain the frame-level anomaly scores, we keep the maximum value from each output map $\boldsymbol{o}_{t}$ and subsequently apply another temporal Gaussian filter to smooth the values.